# Disjoint supercyclicity in Banach algebras

### v7

@@ Abstract.
In this paper, we consider operators that are a composition of an isometric isomorphism and a left multiplier on a normed algebra and we characterize disjoint $\mathcal{F}$-semi-transitive and disjoint supercyclic such operators on a large class of non-unital normed algebras. It turns out that generalized weighted bilateral shifts on the standard Hilbert C*-module is just a special case of our theory. Generalized weighted composition operators on the normed algebra of operator-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space are another special case of our theory. Also, we illustrate our results with concrete examples.
@@
1. Introduction and preliminaries
We recall first the following definition.
@@ 1. Introduction and preliminaries
[18, Definition 4.1] Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
@@
Let now $\mathcal{A}$ be a non-unital Banach algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$. We do not assume that the norm on $\mathcal{A}$ extends to $\mathcal{A}_{1}$. We will say that $\mathcal{A}$ satisfies the condition (E) with respect to $\mathcal{A}_{1}$ if
@@ 2. Main results
For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$
Then $\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
$\text{Similarly, for each }t\in F\text{ and }l\in\{1,\dots,N\},\text{ we have}$
@@
[…]
$\text{ set }d_{t}=a_{t}p_{\alpha}.$
$\text{Then }\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon\quad\text{and}\quad|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
@@
We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system is only needed for the proof of the implication $(1)\Rightarrow(2)$. Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities.
@@
$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}+p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
@@
$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}+p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
@@ 3. Applications
From now on and in the rest of this section we will assume that $\mathcal{F}$ the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
@@
Moreover, we will denote by $\Omega$ be a locally compact Hausdorff, non-compact space. As usual, $C_{b}(\Omega)$ will denote the space of all bounded continuous functions on $\Omega$ and $C_{0}(\Omega)$ will denote the space of all continuous functions on $\Omega$ vanishing at infinity. Both these spaces will be equipped with supremum norm. In addition, we let $C_{c}(\Omega)$ is the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subset $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.
3.1. Applications to generalized weighted shifts on Hilbert C*-modules
In this subsection, we will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [21, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
$\{x_{j}\}_{j\in\mathbb{Z}}\cdot\{y_{j}\}_{j\in\mathbb{Z}}=\{x_{j}y_{j}\}_{j\in\mathbb{Z}}.$
To see that $\{x_{j}y_{j}\}_{j\in\mathbb{Z}}$ belongs to $\ell_{2}(\mathcal{C})$, it suffices to observe that
$\sum_{j}y_{j}^{*}x_{j}^{*}x_{j}y_{j}\leq\sum_{j}y_{j}^{*}\|x_{j}\|^{2}y_{j}\leq\parallel\{x_{j}\}_{j\in\mathbb{Z}}\parallel^{2}\sum y_{j}^{*}y_{j}.$
For each $m,J\in\mathbb{N},$ we let $\tilde{p}_{J,m}\in\ell_{2}(\mathcal{C})$ be given by
[…]
As shown in the proof of [16, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$. We put then $\mathcal{A}=\ell_{2}(\mathcal{C}),$ and
$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in B(H)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|\leq M_{y}\ \forall j\}.$
The multiplication on $\mathcal{A}_{1}$ is defined similarly component-wise. We define the norm $\|\cdot\|_{1}$ on $\mathcal{A}_{1}$ as $\|\{y_{j}\}_{j\in\mathbb{Z}}\|_{1}=sup_{j\in\mathbb{Z}}\|y_{j}\|$ for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1}.$ It is straightforward to check that the condition (E) is satisfied in this case.
Let now $\mathcal{U}$ be a unitary operator on $H$. For each $n\in\mathbb{N}$ we define the map $\Phi_{n}:\mathcal{A}_{1}\to\mathcal{A}_{1}$ by
$\Phi_{n}\left(\{y_{j}\}_{j\in\mathbb{Z}}\right)=\{\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}\}_{j\in\mathbb{Z}},$
for all $\{y_{j}\}_{j\in\mathbb{Z}}\in\mathcal{A}_{1},$ (that is the $j$-th coordinate of $\Phi_{n}(\{y_{j}\}_{j\in\mathbb{Z}})=\mathcal{U}^{-n}y_{j-n}\mathcal{U}^{n}$ for all $j\in\mathbb{Z}$).
Now, the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\mathcal{A}=\ell_{2}(\mathcal{C})$ and $\tilde{p}_{J,m}=\tilde{p}_{J,m}^{3}$ for all $(J,m)\in\mathbb{N}^{2}.$ Moreover, by some calculations it can be checked that $\{\Phi_{n}\}_{n\in\mathbb{N}}$ is aperiodic with respect to the left approximate unit $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ and that the condition (R) is satisfied in this case.
For each $n\in\mathbb{N},$ let $\{W_{n,j}\}_{j\in\mathbb{Z}}$ be a family of operators in $B(H)$ which is uniformly bounded in norm such that each $W_{n,j}$ has a bounded inverse and such that $\{W_{n,j}^{-1}\}_{j\in\mathbb{Z}}$ is also uniformly bounded in the norm. Then, for each $n\in\mathbb{N}$ we let $b_{n}\in\mathcal{A}_{1}$ be given by
$(b_{n})_{i}=W_{n,i}\mathcal{U}^{n},$
for all $i\in\mathbb{Z}.$ Obviously, $b_{n}$ is invertible in $\mathcal{A}_{1}$ for each $n\in\mathbb{N}$ and
$(b_{n}^{-1})_{i}=\mathcal{U}^{-n}\,W^{\,-1}_{n,i},$
for all $i\in\mathbb{Z},$ hence, the condition $(C)$ is satisfied in this case.
Finally, for each $n\in\mathbb{N}$ we let $T_{n}$ be the operator on $\mathcal{A}$ given by
$T_{n}(a)=b_{n}\,\Phi_{n}(a),\qquad\text{for all }a\in\mathcal{A}.$
Noticing that for each $a\in\mathcal{A}$ and $j\in\mathbb{Z}$ it holds that $\|a_{j}\|\leq\|a\|$, in view of Corollary 2.5 and by arguments similar to those in the proofs of [16, Theorem 3.2] and [17, Proposition 2.7] it is not hard to deduce the following proposition in this special case.
Proposition 3.1.
Under the above notation and assumptions, if there exist dense subsets $H_{0}$ and $H_{1}$ of $H$ and a strictly increasing sequence $\{n_{k}\}_{k}$ of natural numbers such that either
$\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W_{n_{k,j+n_{k}}}y\|=\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W_{n_{k},j+n_{k}}y\|$
$\lim_{k\to\infty}\|W_{n_{k},j+2n_{k}}W_{n_{k},j+n_{k}}y\|=0,$
or
$\lim_{k\to\infty}\|W_{n_{k},j+n_{k}}x\|\,\|W^{-1}_{n_{k},j}y\|=\lim_{k\to\infty}\|W^{-1}_{n_{k},j}x\|\,\|W^{-1}_{n_{k},j}y\|=$
$\lim_{k\to\infty}\|W^{-1}_{n_{k},j-n_{k}}W^{-1}_{n_{k},j}y\|=0,$
for all $x\in H_{0}$ and $y\in H_{1}$, then $\{C_{n}\}_{n\in\mathbb{N}}$ is $\mathcal{F}$–semi-transitive on $\mathcal{A}$.
Example 3.2.
Let $H=L^{2}(\mathbb{R})$. As in [16, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel,\parallel w_{j}^{-1}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq K$ for all $j\in\mathbb{Z}$ and some $K>0$. For each $j\in\mathbb{Z},$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R},$ and let then $V_{j}\in B(H)$ be given by
$V_{j}(f)=w_{j}\cdot(f\circ\alpha_{j})$
for all $f\in H.$ For each $n\in\mathbb{N},i\in\mathbb{Z}$, let $W_{n,i}\in B(H)$ be given by
$W_{n,i}:=V_{i}V_{i-1}\cdots V_{i-n+1}.$
If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [16, Example 3.3] one can show that the conditions of Proposition 3.1 are satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R})$ in this case. On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 3.1 will be satisfied with $H_{0}=H_{1}=C_{c}(\mathbb{R}).$
Next, for $N\in\mathbb{N},$ let $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}.$
Let $\alpha_{1},\dots,\alpha_{N}$ be homeomorphisms of $\Omega$, $\{w_{j}^{(1)}\}_{j\in\mathbb{Z}},\dots,\{w_{j}^{(N)}\}_{j\in\mathbb{Z}}\subseteq C_{b}(\Omega)$. We will assume that $w_{j}^{(\ell)}>0$ for all $j\in\mathbb{Z}$ and $\ell\in\{1,\dots,N\}$ and that $\|w_{j}^{(\ell)}\|_{\infty}<M$ for all $j\in\mathbb{Z}$, $\ell\in\{1,\dots,N\}$ and some $M>0.$ Now we consider $\mathcal{A}=\ell_{2}(C_{0}(\Omega))$, and
$\mathcal{A}_{1}:=\{\{y_{j}\}_{j\in\mathbb{Z}}\mid y_{j}\in C_{b}(\Omega)\ \forall j,\,\exists M_{y}>0\text{ such that }\|y_{j}\|_{\infty}\leq M_{y}\ \forall j\}.$
The multiplication and the norm on $\mathcal{A}_{1}$ are defined in the same way as in the previous case. Once again, it is straightforward to check that the condition (E) is satisfied in this case.
For $J\in\mathbb{N}$, we let $\tilde{p}_{(K_{1},K_{2},J)}\in\mathcal{A}$ be given by
[…]
Then it is easily checked that $\{\tilde{p}_{(K_{1},K_{2},J)}^{3}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}$ satisfies the condition $(\mathcal{P}.)$
For each $\ell\in\{1,\dots,N\}$ and $k\in\mathbb{N}$, we let $\tilde{\Phi}_{k,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
$\tilde{\Phi}_{k,\ell}\big(\{f_{j}\}_{j\in\mathbb{Z}}\big)=\{f_{j-k}\circ\alpha_{\ell}^{k}\,\}_{j\in\mathbb{Z}},$
for all $\{f_{j}\}_{j\in\mathbb{Z}}\subset\mathcal{A}$. Let now $r_{1},\dots,r_{N}\in\mathbb{N}$ with $r_{1}<r_{2}<\dots<r_{N}$. For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, set
$\Phi_{n,\ell}=\tilde{\Phi}_{r_{\ell}n,\ell}.$
Clearly, the system of isometric isomorphisms $\{\Phi_{k,1}\}_{k\in\mathbb{N}},\dots,\{\Phi_{k,N}\}_{k\in\mathbb{N}}$ is disjoint aperiodic with respect to
$\{\tilde{p}_{(K_{1},K_{2},J)}\}_{K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact ,}J\in\mathbb{N}}.$
Also, the condition (R) is satisfied in this case.
For each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, let $b_{n,\ell}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell})_{j+r_{\ell}n}=\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)\quad\text{for all }j\in\mathbb{Z}.$
Since for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ the function $w_{j}^{(\ell)}>0,$ it follows that the condition $(C)$ is satisfied in this case.
Now, for each $K_{2}\subseteq\Omega$ compact, we can find some $v_{K_{2}}\in C_{0}(\Omega)$ with $0\leq v_{K_{2}}\leq 1$ such that ${v_{K_{2}}}_{|_{K_{2}}}=1.$ Recalling that the approximate unit for $\mathcal{A}$ in this case is indexed over all triples $(K_{1},K_{2},J)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }J\in\mathbb{N},$ for each $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\}$, and every fixed triple $(K_{1},K_{2},J)$ we let $b_{n,\ell,(K_{1},K_{2},J)}^{-1}\in\mathcal{A}_{1}$ be given by
$(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=\frac{v_{K_{2}}}{\prod_{k=1}^{r_{\ell}n}\big(w_{j+k}^{(\ell)}\circ\alpha_{\ell}^{n-k}\big)}$
for all $j\in\mathbb{Z}$ with $J-r_{\ell}n\leq j\leq J-r_{\ell}n$ and $(b_{n,\ell,(K_{1},K_{2},J)}^{-1})_{j+r_{\ell}n}=0$ else. Then, for all $n\in\mathbb{N}$ and $\ell\in\{1,\dots,N\},$ we let $T_{n,\ell}:\mathcal{A}\to\mathcal{A}$ be given by
$T_{n,\ell}(a)=b_{n,\ell}\,\Phi_{n,\ell}(a)\quad\text{for all }a\in\mathcal{A}.$
The next corollary can be deduced by some calculations from Theorem 2.1.
Corollary 3.3.
Under the above notation and assumptions, the following statements are equivalent:
(1) The sequences of operators
$\{T_{n,1}\}_{n\in\mathbb{N}},\ \dots,\ \{T_{n,N}\}_{n\in\mathbb{N}}$
are $d\mathcal{F}$-semi-transitive on $\mathcal{A}$.
(2) For every $J\in\mathbb{N}$ and each compact subset $K$ of $\Omega$, there exists a strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ such that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha_{s}^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha_{\ell}^{-i}\big)(t)\Bigg)=0$
for all $j\in\mathbb{Z}$ with $J\leq j\leq J$ and every $l,s\in\{1,\dots,N\}$. Moreover, for each distinct $s,\ell\in\{1,\dots,N\}$ and all $j\in\mathbb{Z}$ with $J\leq j\leq J$, it holds that
$\lim_{k\to\infty}\Bigg(\sup_{t\in K}\frac{\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j-r_{\ell}n_{k}-i}^{(\ell)}\circ\alpha_{\ell}^{-i}\circ\alpha_{s}^{r_{s}n_{k}}\big)(t)}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j-r_{s}n_{k}+i}^{(s)}\circ\alpha_{s}^{n_{k}-i}\big)(t)}\Bigg)=0.$
Example 3.4.
Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 3.3 are satisfied in this case.

### v8

# Disjoint $\mathcal{F}$-semi-transitivity in Banach modules

@@ Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [28], in this paper, we consider the concept of disjoint Furstenberg-semi-transitivity for operators that are a composition of an isometric isomorphism and a left multiplier on a normed algebra. Thus, we characterize disjoint $\mathcal{F}$-semi-transitive and disjoint supercyclic such operators on a large class of non-unital normed algebras. Generalized weighted composition operators on the normed algebra of operator-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space are a special case of our theory. At the end of the paper, we characterize disjoint $\mathcal{F}$-semi-transitive adjoints of weighted composition operators acting on the weighted space of Radon measures.
@@
1. Introduction
Supercyclicity is a fundamental concept in the dynamics of operators and numerous papers have been written on the topics related to supercyclicity of operators. For example, Hilden and Wallen in [17] proved that any unilateral backward weighted shift is supercyclic. Afterwards, Salas in [33] characterized supercyclic bilateral weighted shift operators on $l^{p}(\mathbb{Z})$ in terms of a supercyclicity criterion. Supercyclicity of several kinds of operators has also been studied in for instance [12, 28, 30, 36]. Moreover, quite recently, a characterization of supercyclic and topologically semi-transitive weighted composition operators on solid Banach function spaces, in particular Morrey spaces, and Segal algebras have been given in [24]. Further, there is a close connection between supercyclicity and semi-Fredholm theory, see [1, 6, 13]. On the other hand, Furstenberg topological transitivity as a generalization of topological transitivity and hypercyclicity has been studied in several papers, see [2, 5, 15, 16, 27]. Motivated by these facts, in this paper we consider the concept of $\mathcal{F}$-semi-transitivity (Definition 2.1), as a generalization of the concept of topological transitivity for supercyclicity given in [28] and we characterize disjoint $\mathcal{F}$-semi-transitive operators on a large class of non-unital normed algebra (Theorem 3.1 and Corollary 3.2). The operators we study are in fact a composition of an isometric isomorphism and a left multiplier on a non-unital normed algebra. The reason for considering these operators is because many known classes of operators that have earlier been studied in connection with the dynamics, such as bilateral weighted shifts on $l^{2}(\mathbb{Z}),$ weighted composition operators on algebras of functions, generalized weighted shifts on Hilbert C*-modules and so on, turn out to be of this kind.
Next, the dynamics of cosine operator functions has been studied in several papers, see [7, 8, 9, 10, 11, 35, 20, 23, 25, 26, 38]. Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of the published papers, as far as we know. All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the above mentioned operators on normed non-unital algebras, i.e. the operators that are a composition of an isometric isomorphism and a left multiplier. In Proposition 3.4 and Corollary 3.5 we provide some sufficient conditions for the cosine operators functions generated by these operators to be $\mathcal{F}$-semi-transitive on the respective normed algebra.
Now, an attempt to axiomatically study hypercyclicity of operators on C*-algebras has been done in [20, Section 4]. This approach is limited to C*-algebras and does not cover many important and relevant cases in linear dynamics. The approach in Section 3 in this paper is more general because we consider a large class of non-unital normed algebras that are not necessarily C*-algebras. Our approach covers also the case of generalized bilateral weighted shift operators on standard Hilbert C*-modules defined in [19, 20].
In Section 4 we provide a new application in the context of generalized weighted composition operators on the space of operator valued continuous functions. More precisely, we denote by $\mathcal{C}$ the C*-algebra of compact operators on a separable Hilbert space $H,$ and in Corollary 4.2 we characterize disjoint $\mathcal{F}$-semi-transitive generalized weighted composition operators on the algebra $C_{0}(\Omega,\mathcal{C})$ equipped with the supremum-norm, i.e. the normed algebra of all $\mathcal{C}$-valued continuous functions vanishing at infinity on a locally compact, non-compact Hausdorff space $\Omega.$ The generalized weighted composition operators that we study in this context are in fact a composition of a weighted composition operator on $C_{0}(\Omega,\mathcal{C})$ and a right multiplication by a unitary operator on $H.$ Thus, these operators are at a same time a generalization of (ordinary) weighted composition operators on the classical space $C_{0}(\Omega),$ and elementary (wedge) operators on $\mathcal{C}$ considered in [21]. We give also a concrete example where the conditions of Corollary 4.2 are satisfied (Example 4.3).
In Section 5, we study the dynamics of adjoints of weighted composition operators. These adjoints act on the weighted space of Radon measures on a locally compact non-compact Hausdorff space. The motivation for considering the dynamics of operators on weighted spaces of Radon measures comes from the recent paper [22] where the dynamics of weighted composition operators on weighted solid Banach function spaces have been studied. In Theorem 5.5 we characterize the adjoints of weighted composition operators that have a dense set of periodic elements, obtaining in this way an improvement of [23, Proposition 10.3] by deducing equivalent conditions that are weaker than sufficient conditions given in [23, Proposition 10.3] for chaos of the adjoints of weighted composition operators. Next, in Theorem 5.6, we characterize disjoint $\mathcal{F}$-semi-transitive adjoints of weighted composition operators on weighted spaces of Radon measures.
2. Preliminaries
We need first the following definition.
@@ 2. Preliminaries
Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}$ and
$\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$
be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
@@
Definition 2.2.
Let $\mathcal{X}$ be a Banach space, and $(T_{n})_{n\in\mathbb{N}_{0}}$ be a sequence of operators in $B(\mathcal{X})$. A vector $x\in\mathcal{X}$ is called a periodic element of $(T_{n})_{n\in\mathbb{N}_{0}}$ if there exists a constant $N\in\mathbb{N}$ such that for each $k\in\mathbb{N}$, $T_{kN}x=x$. The set of all periodic elements of $(T_{n})_{n\in\mathbb{N}_{0}}$ is denoted by ${\mathcal{P}}((T_{n})_{n\in\mathbb{N}_{0}})$. The sequence $(T_{n})_{n\in\mathbb{N}_{0}}$ is called chaotic if $(T_{n})_{n\in\mathbb{N}_{0}}$ is topologically transitive and ${\mathcal{P}}((T_{n})_{n\in\mathbb{N}_{0}})$ is dense in $\mathcal{X}$}. An operator $T\in B(\mathcal{X})$ is called chaotic if the sequence $\{T^{n}\}_{n\in\mathbb{N}_{0}}$ is chaotic.
Let now $\mathcal{A}$ be a non-unital normed algebra such that $\mathcal{A}$ is a left ideal in a unital normed algebra $\mathcal{A}_{1}$. We do not assume that the norm on $\mathcal{A}$ extends to $\mathcal{A}_{1}$. We will say that $\mathcal{A}$ satisfies the condition (E) with respect to $\mathcal{A}_{1}$ if
@@ 3. Main results
For each $t\in F$ and $r\in\{1,\dots,N\},$ we put $g_{t,r}=\lambda_{t}\,T_{t,r}(a_{t})p_{\alpha}.$ Then
$\|g_{t,r}-p_{\alpha}^{2}\|\;<\;\varepsilon$
and
$\frac{1}{|\lambda_{t}|}\,\bigl\|\Phi_{t,r}^{-1}\!\bigl(b_{t,r,\alpha}^{-1}\bigr)\,\Phi_{t,r}^{-1}(g_{t,r})\bigr)\bigr\|\;=\;\frac{1}{|\lambda_{t}|}\,\|b_{t,r,\alpha}^{-1}g_{t,r}\|\;<\;\varepsilon.$
Similarly, for each $t\in F$ and $l\in\{1,\dots,N\},$ we have
@@
and, as shown above,
$\|a_{t}p_{\alpha}-p_{\alpha}^{2}\|\;<\;\varepsilon$
For each $t\in F$ and $l\in\{1,\dots,N\},$ set $d_{t}=a_{t}p_{\alpha}.$ Then
$\|d_{t}-p_{\alpha}^{2}\|\;<\;\varepsilon$
and
$|\lambda_{t}|\,\bigl\|\Phi_{t,l}^{-1}(b_{t,l})\,d_{t}\bigr\|\;=\;\|\lambda_{t}\,b_{t,l}\,\Phi_{t,l}(a_{t}p_{\alpha})\|\;<\;\varepsilon.$
@@
We notice that the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is a disjoint aperiodic system in Theorem 3.1 is only needed for the proof of the implication $(1)\Rightarrow(2)$. Also, we notice that we have used the condition (R) only in the proof of the implication $(1)\Rightarrow(2)$ to obtain the desired inequalities. Hence, the assumption that $\{\Phi_{t,1}\}_{t\in\mathcal{S}},\cdots,\{\Phi_{t,N}\}_{t\in\mathcal{S}}$ is disjoint aperiodic and the condition (R) are not needed in Corollary 3.2.
@@
$v_{t}:=d_{t}p_{\alpha}x+\frac{2\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}(T_{t}(e_{t}+p_{\alpha}x)+S_{t}(e_{t}p_{\alpha}x)).$
@@
$\frac{\sqrt{\|T_{t}(e_{t}p_{\alpha}y)\|+\|S_{t}(f_{t}p_{\alpha}y)\|}}{\sqrt{\|T_{t}(d_{t}p_{\alpha}x)\|+\|S_{t}(d_{t}p_{\alpha}x)\|}}C_{t}(v_{t})\in\mathcal{V}\text{ for all }t\in F.$
@@ 4. Applications
From now on and in the rest of this section we will assume that $\mathcal{F}$ is the family of all infinite subsets of $\mathbb{N}.$ Further, in this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$.
@@
Moreover, we will denote by $\Omega$ be a locally compact, non-compact Hausdorff, space. As usual, $C_{b}(\Omega)$ will denote the space of all bounded continuous functions on $\Omega$ and $C_{0}(\Omega)$ will denote the space of all continuous functions on $\Omega$ vanishing at infinity. Both these spaces will be equipped with supremum norm. In addition, we let $C_{c}(\Omega)$ be the space of all continuous compactly supported functions on $\Omega.$ For each pair of compact subsets $K_{1}$ and $K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}$ , we let $u_{(K_{1},K_{2})}\in C_{0}(\Omega)$ be such that $u_{{(K_{1},K_{2})}|_{K_{1}}}=1$, $\operatorname{supp}u_{(K_{1},K_{2})}=K_{2}$ and $0\leq u_{(K_{1},K_{2})}\leq 1$.